# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 4 Estimates for convection and reaction terms involving discrete Green’s function
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq CN^{2}\left[{\frac{\varepsilon^{1/4}}{\varepsilon^{2}N^{2}}}+\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\frac{h_{y}\varepsilon^{1/4}}{N^{2}\varepsilon^{3/2}}\right]$
$\;+CN^{2}\varepsilon^{1/4}\left[\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}\varepsilon}\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,y})}\right]$
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{\varepsilon^{2}N^{2}}+\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\frac{\varepsilon^{\frac{3\beta}{2}}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\frac{3\beta}{2}}}{N}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.14)
Consider the first term of the R.H.S. of (4.11) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\frac{C}{N}\quad\quad\quad\quad\quad\quad\quad\text{(By inverse inequality)}$ (4.15)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}$
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
@@
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Lemma 5.3 If the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}$ but $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ (4.16)
Consider the first term of the R.H.S. of (4.16) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IR-R\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,xy})}\log(1/\varepsilon^{3/2})\varepsilon^{3/4}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\varepsilon^{-1/4}}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq\frac{CN}{\varepsilon^{-1/4}}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.17)
Consider the second term of the R.H.S. of (4.16) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq CN[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}\varepsilon^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.18)
Next consider the third term of the R.H.S. of (4.16) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq CN[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,xy})}+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{1/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.19)
Next consider the fourth term of the R.H.S. of (4.16) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq CN\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,xy})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,xy})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{\varepsilon^{2}N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{3/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\left[\frac{\log^{2}(1/\varepsilon)}{N}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N}\right]$
$\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$ (4.20)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}$
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
@@
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Lemma 5.4 When $\alpha>5/2$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq C\frac{\log^{2}({1/\varepsilon^{3/2})}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(b_{1}(Iu-u)_{x},G_{mn})_{\Omega_{c}}$.
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$ (4.21)
Consider the first term of the R.H.S. of (4.21) $(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IR-R\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IR-R\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{c})}\;\varepsilon^{-1/2}\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\quad\alpha>1/2.$ (4.22)
Consider the second term of the R.H.S. of (4.21) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon^{2}N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>5/2.$ (4.23)
Next consider the third term of the R.H.S. of (4.21) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{c})}\;\varepsilon^{-1/2}\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{CN}{\sqrt{\varepsilon}}[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|F_{2,yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq\frac{CN}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon^{1/2}N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon N^{2}}\right]$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>3/2.$ (4.24)
Next consider the fourth term of the R.H.S. of (4.21) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$

### v2

@@ 5 Estimates related to convection and reaction terms
$\leq CN\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,y})}](N\log(1/{\varepsilon}))^{1/2}$
$\leq CN(N\log(1/{\varepsilon}))^{1/2}\left[{\frac{\varepsilon^{1/4}}{\varepsilon^{2}N^{2}}}+\frac{h_{y}h_{x}\varepsilon^{1/4}}{\varepsilon^{3/2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}\varepsilon}\right]\varepsilon^{\alpha}$
$\leq CN(N\log(1/{\varepsilon}))^{1/2}\varepsilon^{1/4}\left[\frac{\varepsilon^{\alpha}}{\varepsilon^{2}N^{2}}+\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N^{2}}\frac{\varepsilon^{\alpha}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\alpha}}{N^{2}}\right]$
$\leq\frac{C\varepsilon^{1/4}(\log(1/{\varepsilon}))^{1/2}}{N^{1/2}}\leq C\frac{(\log(1/{\varepsilon}))^{1/2}}{N}.$ (5.9)
Consider the first term of the R.H.S. of (5.6) is $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Corollary 3.4.1 and furthermore, we obtain
$|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}|$ $\leq\|IR-R\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,y})}\sqrt{|\Omega_{f,y}|}\;\|G_{mn,x}\|_{\Omega_{f,y}}$
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,y})}\log(1/\varepsilon^{3/2})\varepsilon^{1/4}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{C\|G_{mn}\|_{1,\varepsilon}}{\varepsilon^{1/4}}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{f,y})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq\frac{C(N\log(1/{\varepsilon}))^{1/2}}{\varepsilon^{1/4}}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\sqrt{\varepsilon}\log^{2}(1/\varepsilon^{3/2})}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{(\log({1/\varepsilon^{3/2}))^{5/2}}}{N^{3/2}}.$ (5.10)
Hence, when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|\leq C\frac{(\log(1/{\varepsilon}))^{1/2}}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
@@
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|\leq C\frac{(\log(1/{\varepsilon}))^{1/2}}{N}.$
Lemma 5.4
If the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then following estimate hold
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|\leq C\frac{(\log({1/\varepsilon^{3/2}))^{5/2}}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and consider $|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|$ and decompose $u$ as $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|\leq$ $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}|+|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$
$+|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}|+|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}|.$ (5.11)
Consider the first term of the R.H.S. of (5.11) is $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Corollary 3.4.1 and furthermore, we obtain
$|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\|IR-R\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\|IR-R\|_{L^{\infty}(\Omega_{f,xy})}\log(1/\varepsilon^{3/2})\varepsilon^{1/4}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq\frac{C\sqrt{N\log(1/{\varepsilon})}}{\varepsilon^{-1/4}}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon^{3/2}\log^{2}(1/\varepsilon^{3/2})}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{(\log({1/\varepsilon^{3/2}))^{3}}}{N^{3/2}}.$ (5.12)
Consider the second term of the R.H.S. of (5.11) is $|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1 and furthermore, we obtain
$|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\sqrt{N}\log(1/{\varepsilon})\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}\varepsilon^{2}}+\frac{\varepsilon^{3/2}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{N^{2}}\right]$
$\leq C\frac{(\log({1/\varepsilon^{3/2}))^{3}}}{N^{3/2}}.$ (5.13)
Next consider the third term of the R.H.S. of (5.11) is $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\sqrt{N}\log(1/{\varepsilon})\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon^{3/2}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{1/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\frac{(\log({1/\varepsilon^{3/2}))^{3}}}{N^{3/2}}.$ (5.14)
Next consider the fourth term of the R.H.S. of (5.11) is $|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Corollary 3.4.1 and furthermore, we obtain
$|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,xy}}|$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,xy}}\|b_{1}G_{mn,x}\|_{\Omega_{f,xy}}$
@@
$\leq C\varepsilon^{1/4}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,xy})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,xy})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,xy})}](N\log(1/{\varepsilon}))^{1/2}$
$\leq C\sqrt{N\log(1/{\varepsilon})}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{\varepsilon^{2}N^{2}}+\frac{\varepsilon^{3/2}\log^{2}(1/\varepsilon^{3/2})}{\varepsilon^{3/2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon^{3/2})}{\varepsilon N^{2}}\right]$
$\leq C\left[\frac{\log^{2}(1/\varepsilon)}{N^{3/2}}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N^{3/2}}+\frac{\log^{2}(1/\varepsilon^{3/2})}{N^{3/2}}\right]\leq C\frac{(\log({1/\varepsilon^{3/2}))^{3}}}{N^{3/2}}.$ (5.15)
Hence, when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N^{3/2}}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$
@@
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,xy}}|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{5/2}}{N}.$
Lemma 5.5
For $\alpha\geq 5/2$, $\beta\geq 2/3$, the following estimates holds:
  1. (i)
    When the source point $(x_{m},y_{n})\in\Omega_{f,x}\;\text{then}\;|(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}|\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}},$
  2. (ii)
    when the source point $(x_{m},y_{n})\in\Omega_{c}\;\text{then}\penalty 10000\ |(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}|\leq C.$
Proof. (i) Take the source point $(x_{m},y_{n})\in\Omega_{f,x}$, consider $(b_{1}|(Iu-u)_{x},G_{mn})_{\Omega_{c}}|$ and decompose the domain in four parts, we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}|\leq$ $|(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}|+|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}|$
$+$ $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}|+|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}|.$ (5.16)
Consider the first term of the R.H.S. of (5.16) is $|(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}|$. By using inverse inequality, Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|(IR-R,b_{1}G_{mn,x})_{\Omega_{c}}|$ $\leq\|IR-R\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IR-R\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
$\leq CN\|IR-R\|_{L^{\infty}(\Omega_{c})}\|G_{mn}\|_{\Omega_{c}}$
$\leq CN\sqrt{N\log(1/{\varepsilon})}[h_{x}^{2}\|R_{xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|R_{xy}\|_{L^{\infty}(\Omega_{c})}+h_{y}^{2}\|R_{yy}\|_{L^{\infty}(\Omega_{c})}]$
$\leq CN\sqrt{N\log(1/{\varepsilon})}\left[\frac{1}{N^{2}}+\frac{1}{N^{2}}+\frac{1}{N^{2}}\right]\leq\frac{C(\log(1/{\varepsilon}))^{1/2}}{N^{1/2}}.$ (5.17)
Consider the second term of the R.H.S. of (5.16) is $|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{c}}|$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{c})}\sqrt{|\Omega_{c}|}\;\|G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq\frac{C\sqrt{N\log(1/{\varepsilon})}}{\sqrt{\varepsilon}}\left[\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon^{2}N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{\varepsilon N^{2}}+\frac{\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}}{N^{2}}\right]$
$\leq\frac{C\sqrt{\log(1/{\varepsilon})}}{N^{3/2}\varepsilon^{5/2}}\varepsilon^{\alpha}\leq\frac{C(\log(1/{\varepsilon}))^{1/2}}{N^{3/2}}\quad\text{when}\;\alpha\geq 5/2.$ (5.18)
Next consider the third term of the R.H.S. of (5.16) is $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}|$. By using inverse inequality, Theorem 1.2, Theorem 4.1, Lemma 3.4 and furthermore, we obtain
$|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{c}}|$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$
@@
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{c})}\;\|G_{mn}\|_{\Omega_{c}}$
$\leq CN\sqrt{N\log\frac{1}{\varepsilon}}\left[\frac{1}{N^{2}}+\frac{1}{\varepsilon^{1/2}N^{2}}+\frac{1}{\varepsilon N^{2}}\right]\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}$
$\leq\frac{C}{\varepsilon N^{1/2}}\sqrt{\log(1/{\varepsilon})}\varepsilon^{\frac{3\beta}{2}}\leq\frac{C(\log(1/{\varepsilon}))^{1/2}}{N^{1/2}}\quad\text{when}\;\beta\geq 2/3.$ (5.19)
Next consider the fourth term of the R.H.S. of (5.16) is $|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}|$. By using inverse inequality, Theorem 1.2, Theorem 4.1, Lemma 3.3, Lemma 3.4 and furthermore, we obtain
$|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{c}}|$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{c}}\|b_{1}G_{mn,x}\|_{\Omega_{c}}$